# Revisiting the impact of neutrino mass hierarchies on neutrino mass constraints in light of recent DESI data

### v1

@@ Abstract
Recent results from DESI combined with cosmic microwave background data give the tightest constraints on the sum of neutrino masses to date. However, these analyses approximate the neutrino mass hierarchy by three degenerate-mass (DM) neutrinos, instead of the normal (NH) and inverted hierarchies (IH) informed by terrestrial neutrino oscillation experiments. Given the stringency of the upper limits from DESI data, we test explicitly whether the inferred neutrino constraints are robust to the choice of neutrino mass ordering using both Bayesian and frequentist methods. For Planck data alone, we find that the DM hierarchy presents a good approximation to the physically motivated hierarchies while showing a strong dependence on the assumed lower bound of the prior, confirming previous studies. For the combined Planck and DESI baryon acoustic oscillation data, we find that assuming NH ($M_{\mathrm{tot}}<0.13\,\mathrm{eV}$) or IH ($M_{\mathrm{tot}}<0.16\,\mathrm{eV}$) loosens the Bayesian upper limits compared to the DM approximation ($M_{\mathrm{tot}}<0.086\,\mathrm{eV}$). The frequentist analysis shows that the different neutrino models fit the data equally well and the loosening of the constraints can thus be attributed to the lower bounds induced by NH and IH. Overall, we find that the DM hierarchy presents a good approximation to the physically motivated hierarchies also for Planck+DESI data as long as the corresponding lower neutrino mass bounds are imposed.
@@ I Introduction
The Dark Energy Spectroscopic Instrument (DESI) collaboration has recently reported the tightest cosmological neutrino mass bounds to date [1, 2, 3, 4], inferring an upper limit on the sum of neutrino masses of $M_{\mathrm{tot}}<0.082\,\mathrm{eV}$ [^1] Due to an update in the ACT lensing likelihood, the previous tighter constraint, $M_{\mathrm{tot}}<0.072\,\mathrm{eV}$, from Planck+ACT+DESI BAO data reported in [1] loosened to $M_{\mathrm{tot}}<0.082\,\mathrm{eV}$ as pointed out in [5, 6]. [6] when combining DESI baryon acoustic oscillation (BAO) data with cosmic microwave background (CMB) data from the Planck experiment [7, 8] and the Atacama Cosmology Telescope (ACT, [9, 10]). This tightens further when combining these data with the DESI measurement of the full shape of the galaxy power spectrum: $M_{\mathrm{tot}}<0.071\,\mathrm{eV}$ [6]. We quote all upper limits in this paper at $95\%$ confidence limit if not otherwise indicated.
@@
While there is no strong preference for either of the neutrino hierarchies from neutrino oscillation data [13, 14, 15] and upper limits from terrestrial measurements are still weak, e.g. $m_{\nu}\lesssim 0.45\,\mathrm{eV}$ (90% C.L.) for the effective electron-neutrino mass from the KATRIN experiment [16], the tight upper limits from cosmological data, in particular the recent DESI results, seem to be increasingly in favor of the NH (see e.g. [17, 18, 19, 20, 21, 22] for discussions about evidence for either hierarchy in cosmology). This preference for very small, vanishing or even “negative” neutrino masses has been studied extensively in the literature (e.g. [23, 24, 25, 26, 27]).
@@
However, the current DESI results [1, 6] – like most studies – approximate the physically motivated NH and IH by assuming that the three species of massive neutrinos each carry one third of the total mass: $m_{1}=m_{2}=m_{3}=M_{\mathrm{tot}}/3$, which we refer to as the “degenerate-mass hierarchy” (DM). This approximation has been shown [28, 29, 30] to be accurate for WMAP [31] and Planck data. Even optimistic forecasts, including experiments like CMB-S4 [32], LiteBIRD [33], Euclid [34] and SKA [35] seem to be unable to distinguish between the different hierarchies [36]. Ref. [37] finds that using the DM approximation with a lower limit informed from NH or IH gives constraints in good agreement with a full NH or IH analysis in forecasts for the CORE experiment. [^2] Ref. [38] studies non-linear effects induced by the explicit neutrino mass ordering that could help to distinguish these. Moreover, line-intensity mapping surveys could be sensitive to possible neutrino decay channels [39]. Another common approximation is to assume one massive species and two massless ones, $m_{1}=M_{\mathrm{tot}},\ m_{2}=m_{3}=0$, which we denote as “1 massive / 2 massless” (1M). Although commonly used in the literature, the 1M hierarchy has been shown to be an inaccurate approximation of the physically motivated hierarchies for Planck data [28].
@@
While these studies indicate that the DM approximation presents a good approximation for cosmological data, we deem it important to confirm this explicitly for recent DESI BAO data given the stringency of the neutrino mass constraints as well as the possibility that the true neutrino masses may differ from the canonical values used in some of the earlier forecasting work. Hence, the goal of this paper is to check whether the inferred tight upper limits on $M_{\mathrm{tot}}$ are robust to the choice of neutrino mass ordering. To do so, we infer $M_{\mathrm{tot}}$ within both Bayesian methods based on MCMC posteriors and frequentist methods based on profile likelihoods, while considering the DM and 1M approximations and the physically motivated NH and IH. The paper is structured as follows; we describe the data sets and Bayesian/frequentist methodology in Sec. II, we discuss our results in Sec. III and conclude in Sec. IV.
II Data sets and methodology
@@
We consider the full DESI BAO sample [1] (referred to as DESI), which measures the angular diameter distances $D_{A}(z)$ and Hubble distances $D_{H}(z)$ in seven redshift bins from different tracers, taking into account correlations between $D_{A}(z)$ and $D_{H}(z)$ [^3] The likelihood compatible with the Cobaya sampler is publicly available at https://github.com/cosmodesi/desilike. We adapt this likelihood to be compatible with the MontePython sampler. This likelihood and notebooks to reproduce the plots are available at https://github.com/LauraHerold/MontePython_desilike.. We combine the BAO data with CMB data from Planck PR3 TT, TE, EE and lensing power spectra [7, 8] (referred to as Planck). The DESI BAO 2024 baseline results [1] additionally include CMB lensing measurements from Planck PR4 [40] and ACT [10]. However, including these data requires an increase in precision in order to get accurate predictions at non-linear scales, which slows down the computations by several factors. This combined with the increased complexity of the neutrino hierarchy modeling lead to prohibitively slow parameter inference. Moreover, the cosmological constraints from Planck+ACT+DESI show a (small) dependence on the non-linear modeling [5] and the precision settings need to be chosen with care to get accurate enough predictions at small scales [41]. Hence, we exclude this data from our analysis but expect that it will not alter the main conclusions of this work.
@@
We use the Boltzmann solver CLASS [42, 43] for predictions of the linear spectra assuming the $\Lambda$CDM model and halofit for non-linear corrections [44, 45]. We adopt both a Bayesian and frequentist analysis. For the Bayesian analysis, we compute posteriors with the MCMC sampler MontePython [46, 47]. We consider the chains as converged when a Gelman-Rubin criterion of $R-1<0.01$ is reached. For plotting the posteriors, we use getdist [48] allowing for adaptive smoothing.
@@
Since the parameter of interest in this work, $M_{\mathrm{tot}}$, is close to its physical boundary in $M_{\mathrm{tot}}=0$, we make use of the boundary-corrected graphical construction, also known as Feldman-Cousins construction [55, 56] (for a detailed description of this construction in a cosmology context see [49]). This construction relies on the asymptotic assumptions such as normality, Wald’s and Wilks’ theorems [57, 58]. These assumptions have been explored for $M_{\mathrm{tot}}$ within the DM approximation and have been found to be consistent with the asymptotic assumptions for the test case in [49]. Here, we do not explicitly test these assumptions and acknowledge that frequentist coverage might only be approximately given.
III Results
@@
To explore the impact of the neutrino mass ordering on the upper limits on the sum of neutrino masses, $M_{\mathrm{tot}}$, inferred from cosmology, we first revisit constraints from Planck CMB data alone, followed by Planck+DESI data.
III.1 Planck
@@
DM $<0.24\,\mathrm{eV}$ $<0.18\ (0.22/0.26)\,\mathrm{eV}$
1M $<0.37\,\mathrm{eV}$ $<0.20\ (0.26/0.30)\,\mathrm{eV}$
@@
The results of the Bayesian analysis under Planck CMB data are summarized in the left column of Tab. 1 and the posterior corner plot is shown in Fig. 1. For the most commonly used approximation, which assumes three neutrinos with degenerate masses (DM), we find $M_{\mathrm{tot}}<0.24\,\mathrm{eV}$ in agreement with the Planck collaborations’ 2018 result [8]. This constraint loosens to $M_{\mathrm{tot}}<0.37\,\mathrm{eV}$ when adopting the alternative commonly used approximation of one massive neutrino and two massless ones (1M). As pointed out in [28], the 1M approximation is a poor approximation of the physical neutrino mass orderings and should be avoided; here, we consider it regardless for comparison. [^5] Despite its well-known limitations, the 1M approximation is often set as a default configuration in samplers like MontePython and Cobaya. Next, we adopt the physical normal (NH) and inverted hierarchies (IH) with mass splittings informed from neutrino oscillation experiments [11], which lead to a physical lower limit of the sum of neutrino masses (Eq. 1, dotted lines in Fig. 1). As is evident in the bottom right panel of Fig. 1, these lower limits cut off the posterior at higher $M_{\mathrm{tot}}$ leading to loosened constraints compared the DM approximation: $M_{\mathrm{tot}}<0.31\,\mathrm{eV}$ (NH) and $M_{\mathrm{tot}}<0.34\,\mathrm{eV}$ (IH). Hence, for the constraints under Planck data alone, assuming the physically motivated NH and IH leads to a loosening of the constraints compared to the DM assumption, driven by imposing a lower limit, re-confirming the findings in [28, 29, 30]. We explore the impact of the prior further with profile likelihoods below.
@@
With the profile likelihood, we can explore the impact of imposing a lower physical bound $M_{\mathrm{tot}}>0.06\,\mathrm{eV}$ and $M_{\mathrm{tot}}>0.1\,\mathrm{eV}$ motivated by the NH and IH, respectively. Imposing these limits as physical boundaries in the Feldman-Cousins construction loosens the bounds to $M_{\mathrm{tot}}<0.22\,\mathrm{eV}$ and $M_{\mathrm{tot}}<0.26\,\mathrm{eV}$ for the DM approximation. The constraints for the remaining neutrino hierarchies are similar and are quoted in parenthesis in Tab. 1. This allows us to disentangle the effect of the lower limit from the effect of different prediction for the CMB and BAO observables in the different neutrino mass orderings. As can be seen from Tab. 1, he former has a stronger impact on the constraints than the latter.
Extrapolating the profile likelihoods to negative $M_{\mathrm{tot}}$ illustrates the well-known preference of Planck data for zero or even “negative” $M_{\mathrm{tot}}$ [52, 24, 26, 49]. The minima of the extrapolated profile likelihoods lie deep in the un-physical negative regime for all hierarchies, between $-0.6\,\mathrm{eV}$ (IH) and $-0.26\,\mathrm{eV}$ (DM, dashed vertical lines in bottom panel of Fig. 2).
Overall, the frequentist constraints are about $25$-$45\,\%$ tighter than the Bayesian constraints (when comparing NH and IH with matching lower bounds) confirming previous comparisons between Bayesian and frequentist neutrino mass constraint [26, 49]. This in itself is not surprising due to the different interpretations and interval-construction methods in the Bayesian and frequentist frameworks, which can lead to discrepant constraints. These discrepancies can be expected to decrease with better data.
III.2 Planck and DESI

### v3

@@ Abstract
Recent results from DESI combined with cosmic microwave background data give the tightest constraints on the sum of neutrino masses to date. However, these analyses approximate the neutrino mass hierarchy by three degenerate-mass (DM) neutrinos, instead of the normal (NH) and inverted hierarchies (IH) informed by terrestrial neutrino oscillation experiments. Given the stringency of the upper limits from DESI data, we test explicitly whether the inferred neutrino mass constraints are robust to the choice of neutrino mass ordering using both Bayesian and frequentist methods. For Planck data alone, we find that the DM hierarchy presents a good approximation to the physically motivated hierarchies while showing a strong dependence on the assumed lower bound of the prior, confirming previous studies. For the combined Planck and DESI baryon acoustic oscillation data, we find that assuming NH ($M_{\mathrm{tot}}<0.13\,\mathrm{eV}$) or IH ($M_{\mathrm{tot}}<0.16\,\mathrm{eV}$) loosens the Bayesian upper limits compared to the DM approximation ($M_{\mathrm{tot}}<0.086\,\mathrm{eV}$). The frequentist analysis shows that the different neutrino models fit the data equally well and the loosening of the constraints can thus be attributed to the lower bounds induced by NH and IH. Overall, we find that the DM hierarchy presents a good approximation to the physically motivated hierarchies also for Planck+DESI data as long as the corresponding lower neutrino mass bounds are imposed.
@@ I Introduction
The Dark Energy Spectroscopic Instrument (DESI) collaboration has recently reported the tightest cosmological neutrino mass bounds to date [1, 2, 3, 4], inferring an upper limit on the sum of neutrino masses of $M_{\mathrm{tot}}<0.082\,\mathrm{eV}$ [^1] Due to an update in the ACT lensing likelihood, the previous tighter constraint, $M_{\mathrm{tot}}<0.072\,\mathrm{eV}$, from Planck+ACT+DESI BAO data reported in [1] loosened to $M_{\mathrm{tot}}<0.082\,\mathrm{eV}$ as pointed out in [5, 6]. [6] when combining DESI baryon acoustic oscillation (BAO) data with cosmic microwave background (CMB) data from the Planck experiment [7, 8] and the Atacama Cosmology Telescope (ACT, [9, 10]). This tightens further when combining these data with the DESI measurement of the full shape of the galaxy power spectrum, $M_{\mathrm{tot}}<0.071\,\mathrm{eV}$ [6]. [^2] We quote all upper limits in this paper at $95\%$ confidence limit if not otherwise indicated.
@@
While there is no strong preference for either of the neutrino hierarchies from neutrino oscillation data [13, 14, 15] and upper limits from terrestrial measurements are still weak, e.g. $m_{\nu}\lesssim 0.45\,\mathrm{eV}$ (90% C.L.) for the effective electron-neutrino mass from the KATRIN experiment [16], the tight upper limits from cosmological data, in particular the recent DESI results, seem to be increasingly in favor of the NH (see e.g. [17, 18, 19, 20, 21, 22, 23, 24] for discussions about evidence for either hierarchy in cosmology). This preference for very small, vanishing or even “negative” neutrino masses has been studied extensively in the literature (e.g. [25, 26, 27, 28, 29]).
@@
However, the current DESI results [1, 6] – like most studies – approximate the physically motivated NH and IH by assuming that the three species of massive neutrinos each carry one third of the total mass: $m_{1}=m_{2}=m_{3}=M_{\mathrm{tot}}/3$, which we refer to as the “degenerate-mass hierarchy” (DM). This approximation has been shown [30, 31, 17, 32] to be accurate for WMAP [33] and Planck data. Even optimistic forecasts, including experiments like CMB-S4 [34], LiteBIRD [35], Euclid [36] and SKA [37] seem to be unable to distinguish between the different hierarchies [38]. Ref. [39] finds that using the DM approximation with a lower limit informed from NH or IH (Eq. 1) gives constraints in good agreement with a full NH or IH analysis in forecasts for the CORE experiment. [^3] Ref. [40] studies non-linear effects induced by the explicit neutrino mass ordering that could help to distinguish these. Moreover, future weak lensing [41, 23] or line-intensity mapping surveys[42, 43] could be sensitive to possible neutrino decay channels. Another common approximation is to assume one massive species and two massless ones, $m_{1}=M_{\mathrm{tot}},\ m_{2}=m_{3}=0$, which we denote as “1 massive / 2 massless” (1M). Although commonly used in the literature, the 1M hierarchy has been shown to lead to a background evolution and matter power spectrum that are significantly different form NH, IH and DM [44, 30, 45, 46, 47].
@@
While these studies indicate that the DM approximation presents a good approximation for cosmological data, we deem it important to confirm this explicitly for recent DESI BAO data given the stringency of the neutrino mass constraints as well as the possibility that the true neutrino masses may differ from the canonical values used in some of the earlier forecasting work. Hence, the goal of this paper is to check whether the inferred tight upper limits on $M_{\mathrm{tot}}$ are robust to the choice of neutrino mass ordering. To do so, we infer $M_{\mathrm{tot}}$ within both Bayesian methods based on MCMC posteriors and frequentist methods based on profile likelihoods, while considering the DM and 1M approximations and the physically motivated NH and IH. The paper is structured as follows; we give a short review about the effect of massive neutrinos on cosmological observables in Sec. II, we describe the data sets and Bayesian/frequentist methodology in Sec. III, we discuss our results in Sec. IV and conclude in Sec. V.
II Massive neutrino’s impact on cosmology
In this section, we briefly review massive neutrinos’ effects on the geometry and the growth of structure of the universe following [47, 48] (see also [31, 49, 46, 50]), where we focus on the impact on the data sets under consideration in this work (CMB power spectra, CMB lensing and BAO).
The geometry of the universe is affected by the neutrinos’ transition from relativistic to non-relativistic particles at a redshift depending on their total mass, $M_{\mathrm{tot}}$. Even for the most conservative bounds, massive neutrinos become non-relativistic only after recombination, since otherwise the early Integrated Sachs-Wolfe effect would be affected at an unacceptable level [8, 51]. Hence, massive neutrinos contribute to the radiation fraction, $\Omega_{r}$, before recombination, while after recombination, they contribute to the matter fraction, $\Omega_{m}$, on scales larger than the free-streaming scale ($k<k_{\mathrm{fs}}$) and to $\Omega_{r}$ on scales smaller than the free-streaming scale ($k>k_{\mathrm{fs}}$). Since $\Omega_{m}(z)\sim(1+z)^{3}$ redshifts slower than $\Omega_{r}(z)\sim(1+z)^{4}$, this increases the expansion rate, $H(z)=H_{0}\sqrt{\Omega_{r}(1+z)^{4}+\Omega_{m}(1+z)^{3}+\Omega_{\Lambda}}$, in the late universe compared to a universe with massless neutrinos. Taking all other parameters as fixed, this would decrease the angular diameter distance to last scattering, $D_{A}=\int_{0}^{z^{*}}\mathrm{d}z/H(z)$. Since however, the angular size of the sound horizon, $\theta_{s}=r_{s}/D_{A}$, is precisely measured by the CMB [8], where $r_{s}$ is the physical size of the sound horizon, the relative increase in $\Omega_{m}$ needs to be compensated by a lower $H_{0}$, leading to the well-known negative $M_{\nu}$-$H_{0}$-degeneracy. This degeneracy can be partially broken by including BAO data, which is sensitive to both $H(z)$ and $\Omega_{m}(z)$ at a redshift $z$. Thus all additional $\Omega_{m}$ present at the BAO redshift $z$ compared to the CMB-inferred $\Omega_{m}$ can be attributed to massive neutrinos.
Massive neutrinos not only affect the geometry, but also the growth of structure of the universe: The excess energy density of massive neutrinos, and thus larger $H(z)$, leads to a larger Hubble friction, which slows down the growth of perturbations on all scales. On scales larger than $k_{\mathrm{fs}}$, this effect is compensated by the relative increase in $\Omega_{m}$ that leads to increased clustering through gravity. Neutrino free streaming, i.e. the neutrinos’ inability to cluster at $k>k_{\mathrm{fs}}$, suppresses the matter power spectrum at scales smaller than $k_{\mathrm{fs}}$. This suppression of structure growth results in a suppressed CMB lensing signal.
Both effects discussed above are dominated by the total mass, $M_{\mathrm{tot}}$. The distribution of $M_{\mathrm{tot}}$ among the three neutrino species could, in principle, impact cosmological observables, but the effects are small as shown by [38, 52]. We briefly review their results here, comparing the NH/IH scenarios with the DM approximation first, and commenting on 1M in the end. The NH and IH have one ($m_{3}$) or two neutrinos ($m_{1},\,m_{2}$), respectively, that are more massive than the three neutrinos in the DM approximation and thus become non-relativistic earlier. Since matter redshifts slower than radiation, this leads to excess energy density in NH/IH compared to DM at redshifts $z\sim 1-100$. As described above, excess energy leads a decrease in $D_{A}$ and thus to small shifts of $\theta_{s}$, i.e. small (sub-cosmic-variance) shifts of the acoustic peaks in the CMB power spectrum and the BAO scale.
The excess energy density in NH/IH compared to DM impacts the growth of structure similar to the description above: on scales larger than $k_{\mathrm{fs}}$, the increased $H(z)$ leads to a suppressed growth of perturbations. Moreover, the scale $k_{\mathrm{fs}}$ itself experiences small shifts due to the different mass orderings: $k_{\mathrm{fs}}$ is dominated by the mass of the heaviest neutrino, which is larger in NH/IH than in DM. Hence free streaming occurs on smaller scales in NH/IH than in DM. These two effects lead to an overall suppression of CMB lensing in NH/IH compared to DM.
The 1M approximation features one neutrino that is even heavier and becomes non-relativistic even earlier than in the NH/IH scenarios, leading to even more excess energy. Hence, the shifts in $\theta_{s}$ are larger, the growth of perturbations is even more suppressed and the free streaming occurs at even smaller scales than in NH/IH/DM [38].
III Data sets and methodology
@@
We consider the full DESI BAO sample [1] (referred to as DESI), which measures the angular diameter distances $D_{A}(z)/r_{\mathrm{d}}$ and Hubble distances $D_{H}(z)/r_{\mathrm{d}}$ relative to the sound horizon at baryon drag, $r_{\mathrm{d}}$, in seven redshift bins from different tracers, taking into account correlations between $D_{A}(z)/r_{\mathrm{d}}$ and $D_{H}(z)/r_{\mathrm{d}}$. [^4] The likelihood compatible with the Cobaya sampler is publicly available at https://github.com/cosmodesi/desilike. We adapt this likelihood to be compatible with the MontePython sampler. This likelihood and notebooks to reproduce the plots are available at https://github.com/LauraHerold/MontePython_desilike.. We combine the BAO data with CMB data from Planck PR3 TT, TE, EE and lensing power spectra [7, 8] (referred to as Planck), which calibrates above $r_{\mathrm{d}}$. The DESI BAO 2024 baseline results [1] additionally include CMB lensing measurements from Planck PR4 [53] and ACT [10]. However, including these data requires an increase in precision in order to get accurate predictions at non-linear scales, which slows down the computations significantly. This combined with the increased complexity of the neutrino hierarchy modeling led to prohibitively slow parameter inference. Moreover, the cosmological constraints from Planck+ACT+DESI show a (small) dependence on the non-linear modeling [5] and the precision settings need to be chosen with care to get accurate enough predictions at small scales [54]. Hence, we exclude ACT and Planck PR4 data from our analysis but expect that it will not alter the main conclusions of this work.
@@
We use the Boltzmann solver CLASS [55, 56] for predictions of the linear spectra assuming the $\Lambda$CDM model and halofit for non-linear corrections [57, 58]. For all computations in this paper, we fix the number of relativistic degrees of freedom, $N_{\mathrm{eff}}=3.046$. We adopt both a Bayesian and frequentist analysis. For the Bayesian analysis, we compute posteriors with the MCMC sampler MontePython [59, 60]. We sample $M_{\mathrm{tot}}$, while the individual neutrino masses are computed within MontePython using the neutrino oscillation data from [11] and passed to CLASS. We consider the chains as converged when a Gelman-Rubin criterion of $R-1<0.01$ is reached. For plotting the posteriors, we use getdist [61] allowing for adaptive smoothing.
@@
Since the parameter of interest in this work, $M_{\mathrm{tot}}$, is close to its physical boundary in $M_{\mathrm{tot}}=0$, we make use of the boundary-corrected graphical construction, also known as Feldman-Cousins construction [68, 69] (for a detailed description of this construction in a cosmology context see [62]). This construction relies on the asymptotic assumptions such as normality, Wald’s and Wilks’ theorems [70, 71]. These assumptions have been explored for $M_{\mathrm{tot}}$ within the DM approximation and have been found to be consistent with the asymptotic assumptions for Planck-lite data [62]. Here, we do not explicitly test these assumptions and acknowledge that frequentist coverage might only be approximately given.
IV Results
@@
To explore the impact of the neutrino mass ordering on the upper limits in $M_{\mathrm{tot}}$ inferred from cosmology, we first revisit constraints from Planck CMB data alone, followed by Planck+DESI data.
IV.1 Planck
@@
DM $<0.24\,\mathrm{eV}$ $<0.18\ (0.23/0.26)\,\mathrm{eV}$
1M $<0.39\,\mathrm{eV}$ $<0.20\ (0.25/0.29)\,\mathrm{eV}$
@@
The results of the Bayesian analysis under Planck CMB data are summarized in the left column of Tab. 1 and the posterior corner plot is shown in Fig. 1. For the most commonly used approximation, which assumes three neutrinos with degenerate masses (DM), we find $M_{\mathrm{tot}}<0.24\,\mathrm{eV}$ in agreement with the Planck collaborations’ 2018 result [8]. This constraint loosens to $M_{\mathrm{tot}}<0.39\,\mathrm{eV}$ when adopting the alternative commonly used approximation of one massive neutrino and two massless ones (1M). As mentioned above, the 1M approximation leads to significant changes in cosmological observables compared to NH/IH/DM; here, we consider it regardless for comparison. Next, we adopt the physical normal (NH) and inverted hierarchies (IH) with mass splittings informed from neutrino oscillation experiments [11], which lead to a physical lower limit of the sum of neutrino masses (Eq. 1, dotted lines in Fig. 1). As is evident in the bottom right panel of Fig. 1, these lower limits cut off the posterior at higher $M_{\mathrm{tot}}$ leading to loosened constraints compared the DM approximation: $M_{\mathrm{tot}}<0.31\,\mathrm{eV}$ (NH) and $M_{\mathrm{tot}}<0.34\,\mathrm{eV}$ (IH). Hence, for the constraints under Planck data alone, assuming the physically motivated NH and IH leads to a loosening of the constraints compared to the DM assumption, driven by imposing a lower limit, re-confirming the findings in [30, 31, 66, 32]. We explore the impact of the prior further with profile likelihoods below.
@@
With the profile likelihood, we can explore the impact of imposing a lower physical bound $M_{\mathrm{tot}}>0.06\,\mathrm{eV}$ and $M_{\mathrm{tot}}>0.1\,\mathrm{eV}$ motivated by the NH and IH, respectively. Imposing these limits as physical boundaries in the Feldman-Cousins construction loosens the bounds to $M_{\mathrm{tot}}<0.23\,\mathrm{eV}$ and $M_{\mathrm{tot}}<0.26\,\mathrm{eV}$ for the DM approximation. The constraints for the remaining neutrino hierarchies are similar and are quoted in parentheses in Tab. 1. This allows us to disentangle the effect of the lower limit from the effect of different predictions for the CMB and BAO observables in the different neutrino mass orderings. As can be seen from Tab. 1, the former has a stronger impact on the constraints than the latter.
Extrapolating the profile likelihoods to negative $M_{\mathrm{tot}}$ illustrates the well-known preference of Planck data for zero or even “negative” $M_{\mathrm{tot}}$ [65, 26, 28, 62]. The minima of the extrapolated profile likelihoods lie deep in the un-physical negative regime for all hierarchies, between $-3.5\,\mathrm{eV}$ (1M) and $-0.27\,\mathrm{eV}$ (DM, dashed vertical lines in bottom panel of Fig. 2).
Overall, the frequentist constraints are about $25$-$50\,\%$ tighter than the Bayesian constraints (when comparing NH and IH with matching lower bounds) confirming previous comparisons between Bayesian and frequentist neutrino mass constraint [28, 62]. This in itself is not surprising due to the different interpretations and interval-construction methods in the Bayesian and frequentist frameworks, which can lead to discrepant constraints. These discrepancies can be expected to decrease with better data.
IV.2 Planck and DESI